MIDC News
Year of 2015
December 3, 2015
- 10:17 MST: A LICOR LI-1800 spectroradiometer operating on a global 1-axis tracker system was installed at SRRL on December 1. The LI-1800 detector is maintained at 40 degrees C. -Afshin Andreas
- 16:40 MST: The ESR changed their data file format on July 21, 2015, as a result the following post-processed values from the SRRL Prede POM-01 photometer were mislabeled: AOD, AOD-DIF, and SSA (SkyNet/SkyRad Method, Levels 0 and 1.1). As of now, a new data file format has been applied to MIDC to correct this and to show labels for two new values that were added (IDCR and IDCI). Note: For Level 1.1 data this change was applied on July 1, 2015. -Afshin Andreas
- 12:25 MST: The SRRL BMS Upwelling and Downwelling PIRs were swapped out with an instruments calibrated in BORCAL-LW 2015-04 between 11:59 and 12:04 MST today. Coefficients were updated in logger at 12:04 MST today. -Afshin Andreas
- 17:21 MST: The SRRL BMS CG4 was swapped out with an instrument calibrated in BORCAL-LW 2015-04 at 16:29 MST today. Coefficients were updated in logger at 16:30 MST today. -Afshin Andreas
- 10:40 MST: The internal MIDC web server was upgraded from a Dell PowerEdge 1900 Dual Quad Core (8 processor) 2.33 GHz Xeon w/4 GB RAM (installed Jan 2008) to a Dell Precision T7500 Dual Quad Core (8 processor) 3.20 GHz Xeon w/48 GB RAM at 17:45 MST yesterday. -Afshin Andreas
- 11:41 MST: A Global Horizontal LI-200R has been added to SRRL BMS. Data is available since 04/10/2015. -Afshin Andreas
- 14:02 MST: All SRRL BMS Global UV instruments were calibrated on 08-05-2015 this includes: UV-S-B-T, CUVB1, MS210W, UVB-1, 501A, TUVR, CUV4, UV-S-A-T, CUVA1. Logger was updated with new calibration factors today @ 13:56 MST. -Afshin Andreas
- 11:37 MST: A new folder has been created on the MIDC public server to store the graphical display images (because the permissions on the original folder could not be repaired). However, yesterday during troubleshooting the proxy to the midc public server from the main www.nrel.gov server was broken. Until this is resolved external users can access the MIDC public server directly. -Afshin Andreas
- 11:42 MST: The MIDC Public web server was migrated to a new server at 11:09 MST today. However the live graphical display page for all sites is not working due to a permission problem on the folder containing the display images. Internal NREL users can continue to use the internal MIDC server until this is resolved. -Afshin Andreas
- 14:22 MST: The SRRL BMS SE-POA LI-200 was swapped out with a sensor calibrated in BORCAL 2015-03 at 14:06 MST. Calibration factor was updated in logger at 14:19 MST. -Afshin Andreas
- 15:48 MST: The following SRRL BMS Radiometer Tower pyranometers were swapped out with sensors calibrated in BORCAL 2015-03 between 12:40 and 13:05 MST today: PSPs (90-North/East/South/West) and LI-200s (90-North/East/South/West,40-South,Upwelling). Calibration factors were updated in the logger at 13:00 MST. In addition, the offset for the snow depth sensor was updated at the same time. -Afshin Andreas
- 09:35 MST: The temperature, wind speed, and wind direction sensors at NWTC M2 were replaced with calibrated instruments on July 28, 2015 (between 7:30 and 13:50 MST). The data logger program was updated today @ 9:30 MST with new calibration constants for the wind speed and temperature sensors. -Afshin Andreas
- 08:43 MST: The SRRL BMS CR3000s (both deck and rad tower) were calibrated between 13:50 and 15:00 MST yesterday. During that time the following instruments may have a few missing or bogus data points: TSP-700, SP-110, ML-01, LI-201, SB-PSP. -Afshin Andreas
- 16:49 MST: The SRRL BMS Instrument Mounting Locations PDF file has been updated to show changes that occured with BMS version 13. In addition, historical documents since v8 (November 2000) are now available. -Afshin Andreas
- 08:49 MST: The WXT520 was re-installed on the SRRL RSR2 at 08:38 MST today. The WXT520 had been temporarily removed (and replaced with CS106 and CS215) on 09/18/2013 for testing a standard template program. -Afshin Andreas
- 14:43 MST: The filter correction factor for the SRRL BMS Direct NIP RG780 was recomputed and applied at 14:39 MST today using RG780 filter transmission spectra that were measured today. -Afshin Andreas
- 10:07 MST: Photos of SRRL BMS Sci-Tec trackers have been updated. -Afshin Andreas
- 16:21 MST: Some of the SRRL BMS Instrument Pictures (both current and de-commissioned page) have been updated to reflect major BMS upgrades that have occurred in 2015. The following photos still need to be updated: instrumentation on Sci-Tec trackers and instrumentation on radiometer tower. -Afshin Andreas
- 11:44 MST: The Global, Direct, and Diffuse instruments at SolarTAC were swapped out with instruments calibrated in BORCAL 2015-01 between 7:40 and 8:40 MST yesterday. Data logger program was updated with new calibration factors at 8:04 MST yesterday. -Afshin Andreas
- 11:29 MST: The SRRL BMS Upwelling Shortwave CM21 was swapped out with an instrument calibrated in BORCAL 2015-03 at 11:08 MST today. -Afshin Andreas
- 10:49 MST: The OTF MS-700 was calibrated and returned to service at 10:29 MST on July 14. Unit was found in tolerance. -Afshin Andreas
- 09:56 MST: The Loyola Marymount University (LMU) RSR Battery was not taking a full charge since December 2014, as a result there are large gaps in data each day. Battery was replaced July 9, 2015. DATA SHOULD BE USED WITH CAUTION DURING THE PERIOD WITH BATTERY PROBLEMS. -Afshin Andreas
- 13:57 MST: The following EKO thermopile global pyranometers have been added to SRRL BMS: MS-802, MS-410, and MS-602. Data is available since 3/10/2015, instruments were calibrated in BORCAL 2015-01 and new factors were updated in logger on 6/22/2015 at 9:46 MST. -Afshin Andreas
- 16:52 MST: The OTF MS-700 spectroradiometer was removed from service at 8:43 MST today for calibration. -Afshin Andreas
- 14:51 MST: As of 14:42 MST today, the SRRL BMS POM-01 photometer signals (used for real time troubleshooting) are now merged into the BMS data stream by the radiometer tower CR3000 data logger (instead of through post-processing by a PC). The CR3000 pulls the data directly POM-01 Control PC which is running an NREL developed TCP server to provide the POM photometer signal. -Afshin Andreas
- 10:28 MST: The SRRL BMS CR10X logger at the PV/SunEdison Tower was replaced with a CR800 between 8:45 and 9:20 MST. During this time there may be missing data at random times from the following sensors: SE-POA LI-200, SE-POA Angle, WS@22, WS@42, WD@22, and WD@42. At 8:46 MST the HMP-45C Temp/RH probe at the same location was replaced with an HMP60, no data loss occurred for these Temp/RH measurements. -Afshin Andreas
- 16:03 MST: The AM416 multiplexer on the SRRL BMS Radiometer Tower logger was replaced with an AM16/32B between 14:10 and 14:53 MST today. The following instruments will read bogus during this time: LI-200 (40-South, 90-North/East/South/West, and Upwelling), PSP (90-North/East/South/West), LI-190 (Upwelling), GN-Tracker Battery, and EFM. -Afshin Andreas
- 09:55 MST: The NWTC M2 Global PSP was replaced with a calibrated sensor at 09:22 MST today. Calibration factor was updated in data logger program at 09:52 MST. -Afshin Andreas
- 11:56 MST: The SRRL BMS Upwelling Shortwave CM3 was swapped out with an instrument calibrated in BORCAL 2015-02 at 11:54 MST today. -Afshin Andreas
- 10:01 MST: The SRRL BMS Global 40-South CMP11 was swapped out with an instrument calibrated in BORCAL 2015-02 at 10:00 MST today. -Afshin Andreas
- 10:36 MST: SRRL BMS data logger program was updated at 10:32 MST today with a new calibration factor for the Direct DR02 which was calibrated in BORCAL 2015-02. -Afshin Andreas
- 10:35 MST: The SRRL BMS Diffuse 8-48 was swapped out with an instrument calibrated in BORCAL 2015-01 at 10:32 MST today. -Afshin Andreas
- 09:57 MST: A new secondary Direct CHP1 was added to SRRL BMS at 9:49 MST today (calibrated in BORCAL 2015-01). -Afshin Andreas
- 09:49 MST: The existing SRRL BMS Direct CH1 was permanently replaced with a CHP1 at 9:25 MST today. -Afshin Andreas
- 17:37 MST: The SRRL BMS primary Diffuse CM22 was swapped out with an instrument calibrated in BORCAL 2015-01 at 16:45 MST today. -Afshin Andreas
- 17:36 MST: A new secondary Diffuse CM22 was added to SRRL BMS at 16:49 MST today (calibrated in BORCAL 2015-01). -Afshin Andreas
- 15:18 MST: The existing SRRL BMS Global CM22 (Ventilated) was permanently replaced with a CMP22 (Ventilated) at 15:15 MST today. -Afshin Andreas
- 14:27 MST: A new secondary Global CMP22 (unventilated) was added to SRRL BMS at 14:26 MST today (calibrated in BORCAL 2015-01). -Afshin Andreas
- 11:06 MST: The SRRL BMS Diffuse CMP11 was swapped out with an instrument calibrated in BORCAL 2015-01 at 10:54 MST today. -Afshin Andreas
- 10:15 MST: A new Global CMP22 (unventilated) was added to SRRL BMS at 10:03 MST today (calibrated in BORCAL 2015-01). Note: The raw signal (and case temperature) from the instrument has been available since 1-1-2015. -Afshin Andreas
- 10:12 MST: SRRL BMS deck data logger program was updated with new calibration factors for instruments that were calibrated in BORCAL 2015-01 at 9:46 MST today, this includes: CM3, CM6b, TSP-700, CMP11 (global), NIP, GPP, SPP, sNIP, SP-110, SPN1, LI200, LI201SB, ML-01, SP-LITE2, and Global Reserves (1,2,3,4). 2-degree responsivities for PSP (global) were updated at 9:50 MST. -Afshin Andreas
- 07:19 MST: The SRRL BMS Global Normal CM21 was permanently replaced with a CMP11 at 06:53 MST today. -Afshin Andreas
- 09:04 MST: SRRL BMS Custom monthly data is not working on the public MIDC server. NREL employees can use the internal MIDC server. All other users must use the full monthly data sets or the daily data sets. -Afshin Andreas
- 09:33 MST: A DR02 Fast Response Pyrheliometer was added to SRRL BMS at 09:06 MST today. -Afshin Andreas
- 14:42 MST: The anemometer at SolarTAC lost a cup after a hail storm came through on May 15 @ 1:30pm. Wind speed data will be bogus until anemometer is replaced. -Afshin Andreas
- 17:09 MST: As of 16:54 MST today, the SRRL BMS live graphical display as now using data from the new UDP feeds from the Campbell CR3000 data loggers. There is a total of 4 feeds (deck logger @ 3 sec and 1 min; radiometer tower logger @ 3 sec and 1 min). The prior communication method was National Instruments DataSocket and required a PC running LoggerNet for 3 second data and post processing software for the 1 minute data. -Afshin Andreas
- 17:05 MST: As of 5/16/2015 at 07:40 MST, the SRRL BMS cloud cover percentages are now merged into the BMS data stream by the radiometer tower CR3000 data logger (instead of through post-processing by a PC). The CR3000 pulls the data directly from the TSI-880 onboard computer (instead of SQL statements from YESDAQ which was done before). -Afshin Andreas
- 08:24 MST: The SRRL BMS Radiometer Tower CR3000 logger was upgraded to OS 28 today between 7:46 through 7:50 MST. There will be no data from sensors on the radiometer tower during that time. -Afshin Andreas
- 10:41 MST: The MIDC Spectral web interface has been updated to UVA, UVB, and UV-Total Summations. -Afshin Andreas
- 15:18 MST: The Bluefield State College station has been shutdown, last date of data collection is April 13, 2015. -Afshin Andreas
- 13:24 MST: The SMUD Anatolia RSR station has been shutdown, last date of data collection is April 23, 2015. -Afshin Andreas
- 07:35 MST: The SRRL LI-1800 Global Horizontal Spectroradiometer has been discontinued from service as of 06:18 MST today (it has been replaced by the MS-7XX). -Afshin Andreas
- 15:49 MST: The SRRL BMS MS-56 was removed from service at 14:45 MST today, due to possible water contamination which is affecting the instrument signal. Problem started on April 27. -Afshin Andreas
- 08:22 MST: The shadowband on the SRRL TSI-880 Sky Imager failed on April 25, 2015 as a result of the power outage. It will be troubleshot in the future. -Afshin Andreas
- 13:48 MST: The required pressure feed to the SRRL Trimble data acquisition was upgraded to a UDP feed directly from the CR3000 data logger on the radiometer tower at 12:49 MST today. Now PWV data from the Trimble is no longer dependent on a PC running LoggerNet. -Afshin Andreas
- 13:27 MST: Due to a scheduled SRRL power outage, to remove wiring from the building ATS, the following BMS instruments were down on April 25 from 7:20 to 9:39 MST: TSP-700, UVSAT, UVSBT, CUVA1, CUVA2, CUVB1, CUVB2, UVB-1, 501A, MS210W, SPN1, TSI-880, PGS-100, POM-01, TWC, EFM-100, MS-7XX, LI-1800, and Trimble. -Afshin Andreas
- 08:47 MST: There is a scheduled power outage at SRRL today @ 07:20 MST. Duration will be less than 4 hours, MIDC data for all stations may not update during that time. Also certain instrumentation at SRRL BMS will be down during that time. -Afshin Andreas
- 14:48 MST: At 14:46 MST today, the SRRL BMS 40-South PSP was permanently replaced with a CMP11. -Afshin Andreas
- 15:20 MST: At 15:14 MST today the SRRL BMS Upwelling PSP was permanently replaced with a CM21. -Afshin Andreas
- 09:55 MST: The ventilator for the SRRL BMS Diffuse 8-48 was repaired at 09:10 MST today. -Afshin Andreas
- 09:54 MST: The SRRL BMS Diffuse PSP (vent) was permanently removed from service at 08:26 MST today, it will later be replaced by a secondary Diffuse CM22 (vent). -Afshin Andreas
- 12:26 MST: The ventilator fan for the SRRL BMS Diffuse 8-48 has failed. The sunshield was removed at noon to prevent heat build up inside the ventilator housing. -Afshin Andreas
- 15:40 MST: At 15:30 MST today, the SRRL BMS ML-01 sensor was replaced with another sensor (the previous sensor had a leveling problem). -Afshin Andreas
- 08:04 MST: The SolarTAC 2AP Solar Tracker failed at 6:30 MST yesterday. Direct Normal will be zero and Diffuse unshaded. -Afshin Andreas
- 10:46 MST: At 09:14 MST today, an EKO ML-01 silicon sensor was added to SRRL BMS. Also the Apogee SP-110 (new serial number) was returned to service. -Afshin Andreas
- 13:38 MST: At 13:30 MST today, the SRRL BMS Global CM3 (part of CNR1 on Radiometer Tower) was replaced with a stand-alone CM3 radiometer mounted on instrument deck. The CM3 now includes the effective net infrared correction. -Afshin Andreas
- 13:33 MST: At 13:30 MST today, a bug was fixed in SRRL BMS post processing that prevented SBCF from being applied to Shadowband PSP (which was recently re-applied on 2-Jan-2015). -Afshin Andreas
- 12:32 MST: The SRRL BMS Global Quantum LI-190 and Global Photometric sensor LI-210 were relocated from the radiometer tower to the instrument deck between 12:30 and 12:37 MST yesterday. The radiometer tower was a poor location for global measurements as the tower to the north will cause reflections. -Afshin Andreas
- 12:29 MST: A Kipp & Zonen Global CMP11 and Global PQS1 (PAR) sensors were added to SRRL BMS at 15:04 MST yesterday. -Afshin Andreas
- 15:12 MST: The 500nm TWC Photometer was returned to service at 14:50 MST today in SRRL BMS. -Afshin Andreas
- 15:39 MST: The following instruments were added to SRRL BMS recently:
- Data Lab Temp/RH probe was upgraded from HMP-45C to HMP60 on 2-24-2015 @ 16:33 MST
- Global SPP (ventilated) on 2-24-2015 @ 16:33 MST
- Diffuse CMP11 (unventilated) on 2-24-2015 @ 17:40 MST
- Direct sNIP on 2-25-2015 @ 09:53 MST
- Global GPP (unventilated) on 2-25-2015 @ 11:00 MST
-Afshin Andreas - 13:23 MST: Contractors finished signal cabling, junction boxes, and patch panel upgrades at SRRL last week. The temporary BMS data acquisition system which was installed on 11-DEC-2014 was replaced with the permanent system on 24-FEB-2015 @ 16:33 MST. Instruments that were temporarily removed on 11-DEC were returned to service (with the exception of those permanently removed). -Afshin Andreas
- 10:17 MST: SRRL BMS current data format was modified to allow for a future secondary global CMP22 (unventilated) and a future global GPP. Also the Zebra was removed. -Afshin Andreas
- 16:21 MST: January 2014 through December 2014 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 16:18 MST: SRRL BMS Kipp & Zonen SP-Lite was replaced with an SP-Lite2 at 16:10 MST today. Cal factor was updated in logger at 16:12 MST (with value provided by factory). -Afshin Andreas
- 16:28 MST: The following SRRL BMS Radiometer Tower instruments were swapped out with instruments calibrated in BORCAL 2014-02 between 15:05 and 15:45 MST today: Global 90-North/East/South/West PSP & LI-200, Global 40-South PSP & LI-200, Upwelling PSP & LI-200. Calibration factors were updated in the logger in the middle of that time frame. -Afshin Andreas
- 10:16 MST: SRRL BMS data stream was upgraded to version 13 as of 9:01 MST today. Historical data since 1-Jan-2015 has been converted to the new format. However, due signal cable, junction box, and patch panel upgrades taking place this month, no new BMS instrumentation can come online yet (those fields will be filled as missing [-99999] until they go online). However the following changes have occurred now:
- Effective net infrared correction removed from: Global RG780 PSP, Diffuse Shadowband PSP, and Zebra PSP
- Removed from service: Downwelling IR CG3
- SBCF re-applied to Diffuse Shadowband PSP
- Only one Net Radiation reported (instead of two)
- Two Albedos have been computed using different instruments
- Add three Effective Net Infrareds
- Added fields for new instrumentation
- Added more research channels
- 09:39 MST: The ventilator has been permanently removed from the SRRL BMS Global RG780 as of 9:35 MST today. -Afshin Andreas
News Archives:
[1997] [1998] [1999] [2000] [2001] [2002] [2003] [2004] [2005] [2006] [2007] [2008] [2009] [2010] [2011] [2012] [2013] [2014] [2015] [2016] [2017] [2018] [2019] [2020] [2021] [2022] [2023]